OpenStudy (anonymous):

You walk 4 steps north. You turn and then walk 3 steps east. Every step is the same length. What is the shortest distance from where you started to where you ended?

OpenStudy (alexandervonhumboldt2):

pythogorian theorem

OpenStudy (mathmath333):



OpenStudy (alexandervonhumboldt2):

3^2+4^2=x^2 BY PYTH THEOREM

pooja195 (pooja195):

3^2 + 4^2 = c^2 9 + 16 = c^2 25 = c^2 c = √25 Now only figure out the answer

OpenStudy (asnaseer):

Answer depends on where you started from and whether or not you are on a curved or flat surface

OpenStudy (asnaseer):

e.g. consider the case you are on the south pole of the Earth

OpenStudy (anonymous):

you want me to give you guys the choices?

pooja195 (pooja195):

that would be nice

OpenStudy (anonymous):

A.2.6 B.4.9 C.5.0 D.7.0

OpenStudy (asnaseer):

from your choices it looks like the problem is based on a flat surface

OpenStudy (mathmath333):

so use the pythagorus theorem for flat surface that is \(\large \color{black}{\begin{align} x^2=3^2+4^2 \hspace{.33em}\\~\\ \end{align}}\) and solve for \(x\)

OpenStudy (anonymous):

got it thankzz

OpenStudy (mathmath333):

for future reference http://www.mathsisfun.com/pythagorean_triples.html

OpenStudy (mathmath333):

3,4 and 5 is pythagorial triplet